A credit manager rates each applicant for a car loan on a scale of 1 through 5 and then determines the interest rate from the accompanying table FIND THE EQUATION of the linear in slope-intercept form that goes through these points: (1,24)(2,20)(3,16)(4,12)(5,8)
Found 2 solutions by nerdybill, stanbon:
Answer by nerdybill(5404)

.
Slope between two points, can be found by:
m = (y2-y1)/(x2-x1)
.
We can pick any two points but we'll pick 1 and 2 to determine the slope:
m = (y2-y1)/(x2-x1)
m = (20-24)/(2-1)
m = (-4)/(1)
m = -4
.
Now, we can plug any point, say (5,8), and the slope (-4) and plug it into the "point-slope" form:
y-y1 = m(x-x1)
y-8 = -4(x-5)
y-8 = -4x + 20
y = -4x + 28 (this is what they're looking for)



Answer by stanbon(48545)

A credit manager rates each applicant for a car loan on a scale of 1 through 5 and then determines the interest rate from the accompanying table FIND THE EQUATION of the linear in slope-intercept form that goes through these points: (1,24)(2,20)(3,16)(4,12)(5,8)
----------------------------------
slope = (20-24)/(2-1) = -4
------------
intercept: 24 = (-4)(1) = b
b = 28
------------
EQUATION:
y = -4x + 28
===================
Cheers,
Stan H.